# Quantum Dynamics with Stochastic Non-Hermitian Hamiltonians

### v1

@@ Abstract
We study the quantum dynamics generated by a non-Hermitian Hamiltonian subject to stochastic perturbations in its anti-Hermitian part, describing fluctuating gains and losses. The master equation governing the noise-average dynamics describes a new form of dephasing. We characterize the resulting state evolution and analyze its purity. The novel properties of such dynamics are illustrated in a stochastic dissipative qubit. Our analytical results show that adding noise allows for a rich control of the dynamics, with a greater diversity of steady states and the possibility of state purification.
@@
Quantum systems are always in contact with an environment. For this reason, any realistic description of quantum dynamics should include the effects of such coupling. The theory of open quantum systems (OQS) offers a wide range of approaches to this end Breuer and Petruccione (2007); Rivas and Huelga (2012). Two prominent approaches involve non-Hermitian (NH) Ashida et al. (2020); Plenio and Knight (1998); Moiseyev (2011) and stochastic Budini (2001); Chenu et al. (2017); Kiely (2021) Hamiltonians. NH Hamiltonians are generally associated with effective descriptions that account only for a subset of the total number of degrees of freedom of a quantum mechanical system. They can be obtained unraveling the Gorini-Kossakowski-Sudarshan-Lindblad (GKSL) equation Gorini et al. (1976); Lindblad (1976) and post-selecting the trajectories with no quantum jumps Ashida et al. (2020); Roccati et al. (2022), or using projector techniques, as often done in quantum optics, nuclear physics, and molecular quantum chemistry Plenio and Knight (1998); Moiseyev (2011); Feshbach (1958, 1962); Cohen‐Tannoudji et al. (1998). Recently, non-Hermitian Hamiltonians have regained interest because they display novel physical phenomena and topological phases Shen et al. (2018); Gong et al. (2018); Kawabata et al. (2018); Yao and Wang (2018); Takata and Notomi (2018); Yang et al. (2018); Guo et al. (2023) and provide understanding of the quantum Zeno effect Dubey et al. (2023); Snizhko et al. (2020).
@@
NH Hamiltonians generally exhibit complex energy eigenvalues in which the imaginary part describes the lifetime of a state Gamow (1928); Cohen‐Tannoudji et al. (1998). Time evolution ceases then to be trace-preserving due to the leakage of probability density between the subspace under consideration and that associated with the remaining degrees of freedom. The evolution can be made trace-preserving (TP) by renormalizing the density matrix at all times, which renders the evolution nonlinear in the state Carmichael (2008); Brody and Graefe (2012); Alipour et al. (2020); Geller (2023); Rembieliński and Caban (2021). TP nonlinear dynamics has been realized experimentally by post-selection of any dissipative evolution, i.e., discarding the trajectories with quantum jumps Naghiloo et al. (2019); Chen et al. (2021). In this setting, the renormalization of the trace comes naturally as a restriction to the subset of trajectories with no quantum jumps. Furthermore, it allows us to study a plethora of dissipation-induced phenomena Abbasi et al. (2022); Erdamar et al. (2024); Harrington et al. (2022); Chen et al. (2022); Wang et al. (2023).
@@
Stochastic Hamiltonians allow for a treatment of OQS in which the information of the environment is included through the noise statistics and the coupling to the noise Van Kampen (2011); Budini (2001); Kiely (2021). These appear naturally in the stochastic Schrödinger equations obtained by unraveling the GKSL equation in quantum trajectories Wiseman (1996); Wiseman and Milburn (2009); Carmichael (2008). Stochastic Hamiltonians can be used to engineer many-body and long-range interactions Chenu et al. (2017), and their behavior beyond the noise average relates to quantum information scrambling Martinez-Azcona et al. (2023).
@@
Non-Hermitian Hamiltonians are typically studied in a deterministic setting. In this Letter, we go beyond this paradigm by studying the dynamics of NH Hamiltonians subject to classical noise, focusing on the effect of noise in the anti-Hermitian part of the Hamiltonian. Stochastic non-Hermitian Hamiltonians appear, e.g., in homodyne detection of the degenerate parametric oscillator Carmichael (1991). The stability of exceptional points under noise has also been investigated Wiersig (2020). Here, we consider the stochastic dynamics generated by a fluctuating non-Hermitian operator and derive an exact nonlinear master equation for the noise-averaged state dynamics, which recovers noise-induced quantum jumps. In general, this master equation is not of GKSL form and exhibits novel features. We derive an exact equation of motion for the purity and characterize the dynamics. Specifically, we identify the stable steady states in terms of the eigenvalues and eigenstates of the Liouvillian and show that they live in the subspace spanned by the eigenvectors whose eigenvalues’ real part is the largest. We illustrate these regimes by studying the dynamics in a stochastic dissipative qubit Naghiloo et al. (2019). Our analytical results provide the possible steady states, the spectral properties governing the dynamics, and the evolution of the purity.
@@
Dynamics under anti-Hermitian fluctuations. We consider a system with Hermitian Hamiltonian $\hat{H}_{0}$ that is subject to classical noise coupled to the anti-Hermitian operators $i\hat{L}$, i.e.,
@@
This describes fluctuations around the mean in the Hermitian operator $\hat{L}=\hat{L}^{\dagger}$, with standard deviation $\sqrt{2\gamma}$. The classical noise is taken as Gaussian real white noise for simplicity, characterized by $\mathbb{E}(\xi_{t})=0$ and $\mathbb{E}(\xi_{t}\xi_{t^{\prime}})=\delta(t-t^{\prime})$, where $\mathbb{E}(\bullet)$ denotes the classical stochastic average. The (unnormalized) system density matrix for a single trajectory $\tilde{\varrho}_{t}$ evolves, over a small time increment, as $\tilde{\varrho}_{t+\mathrm{d}t}=\hat{U}_{\mathrm{d}t}\tilde{\varrho}_{t}\hat{U}^{\dagger}_{\mathrm{d}t}$, where the propagator $\hat{U}_{\mathrm{d}t}=\exp(-i\hat{H}_{0}\mathrm{d}t-\hat{L}\mathrm{d}t-\sqrt{2\gamma}\hat{L}\mathrm{d}W_{t})$ depends on the Wiener process $\mathrm{d}W_{t}=\xi_{t}\mathrm{d}t$. For Gaussian white noise, the latter obeys Itō’s rules Gardiner (1985): $\mathrm{d}W_{t}^{2}=\mathrm{d}t$ and $(\mathrm{d}W_{t})^{n}=0,\;\forall\;n>2$. Expanding the exponential dictating the evolution of the density matrix with these rules, we find a stochastic differential equation (SDE) for its evolution,
$\mathrm{d}\tilde{\varrho}_{t}=$ $\left(-i[\hat{H}_{0},\tilde{\varrho}_{t}]-\{\hat{L},\tilde{\varrho}_{t}\}+\gamma\{\hat{L},\{\hat{L},\tilde{\varrho}_{t}\}\}\right)\mathrm{d}t$
@@
We denote $\tilde{\rho}\equiv\mathbb{E}(\tilde{\varrho})$ the noise-averaged density matrix and use the tilde $\tilde{\bullet}$ to highlight that it is not normalized. Trace preservation can be imposed either by (i) renormalizing all single-trajectories and then average over the noise Carmichael (2008); Wiseman and Milburn (2009); Dubey et al. (2023); Benoist et al. (2021), or (ii) first averaging over the noise, and only then imposing trace preservation. These two approaches will generally yield different dynamics. We focus on the latter to capture the evolution of ensembles of trajectories and defer the former for future work.
@@
$\frac{\mathrm{d}\tilde{\rho}_{t}}{\mathrm{d}t}=-i[\hat{H}_{0},\tilde{\rho}_{t}]-\{\hat{L},\tilde{\rho}_{t}\}+\gamma\{\hat{L},\{\hat{L},\tilde{\rho}_{t}\}\}\equiv\tilde{\mathcal{L}}[\tilde{\rho}_{t}]$ (3)
for the noise-averaged density matrix, where $\tilde{\mathcal{L}}[\bullet]$ denotes the non-trace-preserving Liouvillian superoperator. This master equation can be formally solved as $\tilde{\rho}_{t}=e^{\tilde{\mathcal{L}}t}[\hat{\rho}_{0}]$. Imposing trace preservation at the average level yields the nonlinear master equation
$\frac{\mathrm{d}\hat{\rho}_{t}}{\mathrm{d}t}=$$\;\tilde{\mathcal{L}}[\hat{\rho}_{t}]-\mathrm{Tr}(\tilde{\mathcal{L}}[\hat{\rho}_{t}])\hat{\rho}_{t},$ (4)
$=$ $-i[\hat{H}_{0},\hat{\rho}_{t}]-\{\hat{L},\hat{\rho}_{t}\}+\gamma\{\hat{L},\{\hat{L},\hat{\rho}_{t}\}\}$
@@
the solution of which is simply given by normalizing the non-trace-preserving (nTP) evolution, i.e., $\hat{\rho}_{t}\equiv\frac{\tilde{\rho}_{t}}{\mathrm{Tr}(\tilde{\rho}_{t})}=e^{\tilde{\mathcal{L}}t}[\hat{\rho}_{0}]/\mathrm{Tr}(e^{\tilde{\mathcal{L}}t}[\hat{\rho}_{0}]).$ Since the NH Hamiltonian is stochastic, the evolution leads to quantum jumps—terms of the form $\gamma\hat{L}\hat{\rho}_{t}\hat{L}$. Note that this master equation is not of the GKSL form; rather, it describes a new form of dephasing in which the jump operators act on the density matrix through a double anticommutator and do not conserve the norm of the state. A more general form is detailed in App. F. Similar equations, with a double anticommutator but with a minus sign, appear when studying the adjoint Liouvillian with fermionic jump operators and evolving operator Liu et al. (2023); Clark et al. (2010).
Note that, in general, the unnormalized evolution (3) can be trace-decreasing (TD), trace-preserving (TP), or trace-increasing (TI). While TD dynamics can easily be interpreted through post-selection, TI evolutions are conceptually challenging, even if sometimes deemed non-physical Nielsen and Chuang (2010) they can describe a biased ensemble of trajectories Garrahan and Lesanovsky (2010). Note that the present TI dynamics can always be made TD or TP by a gauge transformation in the master equation, namely, using an imaginary offset $-ia$ in (1) (cf. App. A).
@@
Purity dynamics. The purity $P_{t}=\mathrm{Tr}(\hat{\rho}_{t}^{2})$ evolves according to the differential equation
@@
Long time dynamics: Stable steady states. To characterize the long-time dynamics, we use the right and left eigendecompositions of the non-trace-preserving Liouvillian superoperator, ~$\tilde{\mathcal{L}}[\tilde{\rho}_{\nu}]=\lambda_{\nu}\tilde{\rho}_{\nu}$ and ~$\tilde{\mathcal{L}}^{\dagger}[\tilde{\rho}_{\nu}^{\textsc{(l)}}]=\lambda_{\nu}^{*}\tilde{\rho}_{\nu}^{\textsc{(l)}}$, respectively. We define the eigenoperator basis as $\hat{\sigma}_{\nu}=\hat{\rho}_{\nu}/\mathrm{Tr}(\hat{\rho}_{\nu})$ if the operator has non-zero trace and as $\hat{\sigma}_{\nu}=\hat{\rho}_{\nu}$ if the operator is traceless. Furthermore, the eigenstate is physical when it is trace one, $\mathrm{Tr}(\sigma_{\nu})=1$, and positive semidefinite $\hat{\sigma}_{\nu}\geq 0$. The basis is ordered such that the physical states appear first $\nu<N_{\textsc{p}}$, the traceful states appear second $\nu<N_{\textsc{t}}$, and the traceless states come last. In each of the subsets, the states are further ordered by decreasing $\mathrm{Re}(\lambda_{\nu})$. Assuming the Liouvillian $\tilde{\mathcal{L}}$ to be diagonalizable, the evolution of the density matrix in this operator basis follows as
@@
At long times, the stable steady state $\hat{\sigma}_{0}$ is the eigenvector associated with the eigenvalue with the largest real part over which there is an initial support, i.e., $\mathrm{Re}(\lambda_{0})=\max_{\nu}\mathrm{Re}(\lambda_{\nu})$ such that $c_{0}\neq 0$. The corrections to this state are suppressed at a rate dictated by the dissipative gap $\Delta=\mathrm{Re}(\lambda_{0}-\lambda_{1})$. They can oscillate with a frequency $\omega_{\nu}=\mathrm{Im}(\lambda_{\nu})$, depending on the presence or absence of complex eigenvalues in the set of states with the second largest real part, denoted by $\mathbb{M}_{2}$.
@@
The long-time limit above assumes that the largest eigenvalue is real and non-degenerate. In a more general case, denoting $\mathbb{M}_{1}$ the set of states with the largest real part among those on which the initial state has nonzero support, the steady state follows as $\hat{\rho}^{\textsc{s}}=\frac{\sum_{\nu\in\mathbb{M}_{1}}^{N^{2}-1}c_{\nu}e^{i\omega_{\nu}t}\hat{\sigma}_{\nu}}{\sum_{\mu\in\mathbb{M}_{1}}^{N_{\textsc{t}}-1}c_{\mu}e^{i\omega_{\mu}t}}$.
@@
Illustration: The stochastic dissipative qubit. Many of the exciting phenomena displayed by non-Hermitian Hamiltonians can be experimentally observed in a dissipative qubit Naghiloo et al. (2019); Chen et al. (2021); Abbasi et al. (2022); Erdamar et al. (2024); Harrington et al. (2022); Chen et al. (2022); Wang et al. (2023). The setting corresponds to the Lindblad evolution of a 3-level system {|$\{\ket{g},\ket{e},\ket{f}\}$, where all quantum jumps to |$\ket{g}$ are discarded. This post-selection process leads to the effectively non-Hermitian Hamiltonian $\hat{H}=J\hat{\sigma}_{x}-i\Gamma\hat{\Pi}$, where $J$ is the hopping between |$\ket{e}$ and |$\ket{f}$, $\Gamma$ is the loss rate from |$\ket{e}$ to |$\ket{g}$ and $\hat{\Pi}=\ket{e}\bra{e}$ | denotes the projector over state |$\ket{e}$. The (shifted) spectrum $\varepsilon_{\pm}+i\frac{\Gamma}{2}=\pm\sqrt{J^{2}-\frac{\Gamma^{2}}{4}}$ is real for $J\geq\frac{\Gamma}{2}$, and imaginary for $J<\frac{\Gamma}{2}$. These two regimes respectively correspond to (passive) $\mathcal{PT}$ unbroken and broken symmetry phases Bender and Boettcher (1998).
@@
Figure 1: Purity evolution for different strengths of the noise as given by (9) in a cross-section of the Bloch sphere.
@@
Figure 2: Spectral and steady-state phase diagrams: (a) dissipative gap $\Delta$, (b) maximum imaginary part of the eigenvalues $\omega_{\nu}$, (c) $z$, and (d) $y$ components of the Bloch vector for the stable steady state. The relevant parameters are the strength of the noise $\gamma$ and the decay rate $\Gamma$. The transition to the noise-induced phase is at $\gamma^{*}=1/(2\Gamma)$ (black). The transition from $\mathcal{PT}$ broken to the $\mathcal{PT}$ unbroken phase happens at $\Gamma/J=2$ (gray dashed). In (b) the power laws $\gamma=1/(3\Gamma)$ and $\gamma=1/\Gamma$ are also shown (orange dotted).
@@
which physically corresponds to considering Gaussian fluctuations in the decay parameter $\Gamma$, with strength $\gamma$. The master equation (4) for this system reads

### v2

@@ Abstract
We study the quantum dynamics generated by a non-Hermitian Hamiltonian subject to stochastic perturbations in its anti-Hermitian part, describing fluctuating gains and losses. The dynamics averaged over the noise is described by an ‘anti-dephasing’ master equation. We characterize the resulting state evolution and analyze its purity. The properties of such dynamics are illustrated in a stochastic dissipative qubit. Our analytical results show that adding noise allows for a rich control of the dynamics, stabilizing the lossy state and making state purification possible to a greater variety of steady states.
@@
Introduction— Quantum systems are always in contact with an environment. For this reason, any realistic description of quantum dynamics should include the effects of such coupling. The theory of open quantum systems (OQS) offers a wide range of approaches to this end Breuer and Petruccione (2007); Rivas and Huelga (2012). Two prominent approaches involve non-Hermitian (NH) Ashida et al. (2020); Plenio and Knight (1998); Moiseyev (2011) and stochastic Budini (2001); Chenu et al. (2017); Kiely (2021) Hamiltonians. NH Hamiltonians are generally associated with effective descriptions that account only for a subset of the total number of degrees of freedom of a quantum mechanical system. They can be obtained by unraveling the Gorini-Kossakowski-Sudarshan-Lindblad (GKSL) equation Gorini et al. (1976); Lindblad (1976) and post-selecting the trajectories with no quantum jumps Ashida et al. (2020); Roccati et al. (2022), or using projector techniques, as often done in quantum optics, nuclear physics, and molecular quantum chemistry Plenio and Knight (1998); Moiseyev (2011); Feshbach (1958, 1962); Cohen‐Tannoudji et al. (1998). Recently, non-Hermitian Hamiltonians have regained interest because of the novel physical phenomena and topological phases they display Shen et al. (2018); Gong et al. (2018); Kawabata et al. (2018); Yao and Wang (2018); Takata and Notomi (2018); Yang et al. (2018); Guo et al. (2023), and of their relation to the quantum Zeno effect Allcock (1969); Schulman (1998); Streed et al. (2006); Echanobe et al. (2008); Dubey et al. (2023); Snizhko et al. (2020).
@@
NH Hamiltonians generally exhibit complex energy eigenvalues in which the imaginary part describes the lifetime of a state Gamow (1928); Cohen‐Tannoudji et al. (1998). Time evolution ceases to be trace-preserving due to the leakage of probability density between the subspace under consideration and that associated with the remaining degrees of freedom. The evolution is made trace-preserving (TP) by renormalizing the density matrix at all times, which renders the evolution nonlinear in the state Carmichael (2008); Brody and Graefe (2012); Alipour et al. (2020); Geller (2023); Rembieliński and Caban (2021). TP nonlinear dynamics has been realized experimentally by post-selection of dissipative evolution, i.e., discarding the trajectories with quantum jumps Naghiloo et al. (2019); Chen et al. (2021). In this setting, the renormalization of the trace comes naturally as a restriction to the subset of trajectories with no quantum jumps. This allowed studying a plethora of dissipation-induced phenomena Abbasi et al. (2022); Erdamar et al. (2024); Harrington et al. (2022); Chen et al. (2022); Wang et al. (2023).
@@
Stochastic Hamiltonians allow for a treatment of OQS in which the effect of the environment is modeled through the noise statistics Van Kampen (2011); Budini (2001); Kiely (2021). These appear naturally in the stochastic Schrödinger equations obtained by unraveling the GKSL equation in quantum trajectories Wiseman (1996); Wiseman and Milburn (2009); Carmichael (2008). Stochastic Hamiltonians have been used to engineer many-body and long-range interactions Chenu et al. (2017), and their behavior beyond the noise average relates to quantum information scrambling Martinez-Azcona et al. (2023).
@@
While the stability of certain features of NH Hamiltonians under noise has recently been investigated Wiersig (2020), non-Hermitian Hamiltonians are typically studied in a deterministic setting. In this Letter, we go beyond this paradigm by studying the dynamics of NH Hamiltonians subject to classical noise. This is particularly relevant since current experimental realizations of NH dynamics use Noisy Intermediate Scale Quantum (NISQ) devices Preskill (2018). We choose to focus on the effect of fluctuations on the anti-Hermitian part, which arise from noise in experimental setups Naghiloo et al. (2019) or in homodyne detection of the degenerate parametric oscillator Carmichael (1991), the possible experimental setups are detailed in App. B. By considering the stochastic dynamics of a fluctuating non-Hermitian system, we unveil a new kind of averaged master equation, that we call ‘anti-dephasing’ because of the structure of the dissipator—which involves anticommutators instead of the commutators commonly encountered in the GKSL dephasing master equation. We derive an exact evolution equation for the purity and characterize its dynamics. Specifically, we identify the stable steady states in terms of the eigenvalues and eigenstates of the Liouvillian. We illustrate the anti-dephasing dynamics in an experimentally-feasible stochastic dissipative qubit Naghiloo et al. (2019); Quinn et al. (2023) (cf. App. B), in which case we find a novel noise-induced phase, where the noise stabilizes the originally lossy state.
@@
Dynamics under anti-Hermitian fluctuations— We consider a system with Hermitian Hamiltonian $\hat{H}_{0}$ that is subject to classical noise coupled to the anti-Hermitian operator $i\hat{L}$, i.e.,
@@
This describes fluctuations around the mean in the Hermitian $\hat{L}=\hat{L}^{\dagger}$, positive $\hat{L}\geq 0$ operator, with standard deviation $\sqrt{2\gamma}$, and could be amenable to experimental implementation (cf. App. B). The classical noise is taken as Gaussian real white noise for simplicity, characterized by $\mathbb{E}(\xi_{t})=0$ and $\mathbb{E}(\xi_{t}\xi_{t^{\prime}})=\delta(t-t^{\prime})$, where $\mathbb{E}(\bullet)$ denotes the classical stochastic average. The (unnormalized) system density matrix for a single trajectory $\tilde{\varrho}_{t}$ evolves, over a small time increment, as $\tilde{\varrho}_{t+\mathrm{d}t}=\hat{U}_{\mathrm{d}t}\tilde{\varrho}_{t}\hat{U}^{\dagger}_{\mathrm{d}t}$, where the propagator $\hat{U}_{\mathrm{d}t}=\exp(-i\hat{H}_{0}\mathrm{d}t-\hat{L}\mathrm{d}t-\sqrt{2\gamma}\hat{L}\mathrm{d}W_{t})$ depends on the Wiener process $\mathrm{d}W_{t}=\xi_{t}\mathrm{d}t$. For Gaussian white noise, the latter obeys Itō’s rules Gardiner (1985): $\mathrm{d}W_{t}^{2}=\mathrm{d}t$ and $(\mathrm{d}W_{t})^{n}=0,\;\forall\;n>2$. Expanding the exponential dictating the evolution of the density matrix with these rules, we find a stochastic differential equation (SDE) for its evolution,
$\mathrm{d}\tilde{\varrho}_{t}=$ $\left(-i[\hat{H}_{0},\tilde{\varrho}_{t}]-\{\hat{L},\tilde{\varrho}_{t}\}+\gamma\big{\{}\hat{L},\{\hat{L},\tilde{\varrho}_{t}\}\big{\}}\right)\mathrm{d}t$
@@
We denote $\tilde{\rho}\equiv\mathbb{E}(\tilde{\varrho})$ the noise-averaged density matrix and use the tilde $\tilde{\bullet}$ to highlight that it is not normalized. Trace preservation can be imposed either by (i) renormalizing all single-trajectories and then averaging over the noise Carmichael (2008); Wiseman and Milburn (2009); Dubey et al. (2023); Benoist et al. (2021), or (ii) first averaging over the noise, and only then imposing trace preservation. These two approaches will generally yield different dynamics. We focus on the latter to capture the evolution of ensembles of trajectories analytically and defer the former for future analysis.
@@
$\frac{\mathrm{d}\tilde{\rho}_{t}}{\mathrm{d}t}=-i[\hat{H}_{0},\tilde{\rho}_{t}]-\{\hat{L},\tilde{\rho}_{t}\}+\gamma\big{\{}\hat{L},\{\hat{L},\tilde{\rho}_{t}\}\big{\}}\equiv\tilde{\mathcal{L}}[\tilde{\rho}_{t}],$ (3)
for the noise-averaged density matrix, where $\tilde{\mathcal{L}}[\bullet]$ denotes the Liouvillian superoperator, which is not necessarily trace-preserving. This master equation can be formally solved as $\tilde{\rho}_{t}=e^{\tilde{\mathcal{L}}t}[\hat{\rho}_{0}]$. Imposing trace preservation at the average level yields the nonlinear master equation
$\frac{\mathrm{d}\hat{\rho}_{t}}{\mathrm{d}t}=$$\;\tilde{\mathcal{L}}[\hat{\rho}_{t}]-\mathrm{Tr}(\tilde{\mathcal{L}}[\hat{\rho}_{t}])\hat{\rho}_{t}$ (4)
$=$ $-i[\hat{H}_{0},\hat{\rho}_{t}]-\{\hat{L},\hat{\rho}_{t}\}+\gamma\big{\{}\hat{L},\{\hat{L},\hat{\rho}_{t}\}\big{\}}$
@@
the solution of which is given by normalizing the non-trace-preserving (nTP) evolution, i.e., $\hat{\rho}_{t}\equiv\frac{\tilde{\rho}_{t}}{\mathrm{Tr}(\tilde{\rho}_{t})}=e^{\tilde{\mathcal{L}}t}[\hat{\rho}_{0}]/\mathrm{Tr}(e^{\tilde{\mathcal{L}}t}[\hat{\rho}_{0}]).$ Since the NH Hamiltonian is stochastic, the evolution leads to quantum jumps—terms of the form $\gamma\hat{L}\hat{\rho}_{t}\hat{L}$. Note that this master equation is not of GKSL form; instead, the jump operators act on the density matrix through a double anticommutator and do not conserve the norm of the state. We term this new dissipator ‘anti-dephasing’. A more standard form is detailed in App. H. Similar equations, with a double anticommutator but with a minus sign, appear when studying the adjoint Liouvillian with fermionic jumps and evolving operators Liu et al. (2023); Clark et al. (2010).
Note that, in general, the unnormalized evolution (3) can be trace-decreasing (TD), trace-preserving (TP), or trace-increasing (TI). While TD dynamics can easily be interpreted through post-selection, e.g., with hybrid Liouvillians Minganti et al. (2020, 2019); Gupta et al. (2024), TI evolutions are conceptually challenging; even if sometimes deemed non-physical Nielsen and Chuang (2010) they can describe a biased ensemble of trajectories Garrahan and Lesanovsky (2010); Lebowitz and Spohn (1999); Esposito et al. (2009); Carollo et al. (2018, 2021); Cech et al. (2023). Remarkably, the present TI dynamics can always be made TD or TP by a gauge transformation in the master equation, namely, using an imaginary offset $-ia$ in (1) (cf. App. A). This is because such a transformation rigidly shifts the Liouvillian spectrum, conserving the relative differences between eigenvalues and, hence, the dynamical features.
@@
Purity dynamics— The purity $P_{t}=\mathrm{Tr}(\hat{\rho}_{t}^{2})$ quantifies how mixed a state is, even when its trace evolves in time. For anti-dephasing dynamics, it evolves according to the differential equation
@@
Long time dynamics: Stable steady states— To characterize the long-time dynamics, we use the right and left eigendecompositions of the Liouvillian superoperator, ~$\tilde{\mathcal{L}}[\tilde{\rho}_{\nu}]=\lambda_{\nu}\tilde{\rho}_{\nu}$ and ~$\tilde{\mathcal{L}}^{\dagger}[\tilde{\rho}_{\nu}^{\textsc{(l)}}]=\lambda_{\nu}^{*}\tilde{\rho}_{\nu}^{\textsc{(l)}}$, respectively. We define the eigenoperator basis as $\hat{\sigma}_{\nu}=\tilde{\rho}_{\nu}/\mathrm{Tr}(\tilde{\rho}_{\nu})$ if the operator has nonzero trace and as $\hat{\sigma}_{\nu}=\tilde{\rho}_{\nu}$ if the operator is traceless. Furthermore, the eigenstate is physical when it has unit trace, $\mathrm{Tr}(\hat{\sigma}_{\nu})=1$, and is positive semidefinite $\hat{\sigma}_{\nu}\geq 0$. The basis is ordered such that the physical states appear first $0\leq\nu<N_{\textsc{p}}$, the traceful states appear second $N_{\textsc{p}}\leq\nu<N_{\textsc{t}}$, and the traceless states come last. In each subset, the states are further ordered by decreasing $\mathrm{Re}(\lambda_{\nu})$. Assuming the Liouvillian $\tilde{\mathcal{L}}$ to be diagonalizable, the evolution of the density matrix in this operator basis follows as
@@
At long times, the stable steady state $\hat{\sigma}_{0}$ is the eigenoperator associated with the eigenvalue with the largest real part over which there is initial support, i.e., $\mathrm{Re}(\lambda_{0})=\max_{\nu}\mathrm{Re}(\lambda_{\nu})$ such that $c_{\nu}\neq 0$. The corrections to this state are suppressed at a rate dictated by the dissipative gap $\Delta=\mathrm{Re}(\lambda_{0}-\lambda_{1})$. They can oscillate with a frequency $\omega_{\nu}=\mathrm{Im}(\lambda_{\nu})$, depending on the presence or absence of complex eigenvalues in the set of states with the second largest real part, denoted by $\mathbb{M}_{2}$.
@@
The long-time limit above (6) assumes that the largest eigenvalue is real and non-degenerate. In a more general case, denoting by $\mathbb{M}_{1}$ the set of states with the largest real part among those on which the initial state has nonzero support, the steady state follows as $\hat{\rho}^{\textsc{s}}=\frac{\sum_{\nu\in\mathbb{M}_{1}}^{N^{2}-1}c_{\nu}e^{i\omega_{\nu}t}\hat{\sigma}_{\nu}}{\sum_{\mu\in\mathbb{M}_{1}}^{N_{\textsc{t}}-1}c_{\mu}e^{i\omega_{\mu}t}}$.
@@
Illustration: The stochastic dissipative qubit— Many of the exciting phenomena displayed by non-Hermitian Hamiltonians have been experimentally observed in a dissipative qubit Naghiloo et al. (2019); Chen et al. (2021); Abbasi et al. (2022); Erdamar et al. (2024); Harrington et al. (2022); Chen et al. (2022); Wang et al. (2023). The setting corresponds to the Lindblad evolution of a 3-level system {|$\{\ket{g},\ket{e},\ket{f}\}$, where all quantum jumps from |$\ket{e}$ to |$\ket{g}$ are discarded. This post-selection process leads to the non-Hermitian Hamiltonian $\hat{H}=J\hat{\sigma}_{x}-i\Gamma\hat{\Pi}$, where $J$ is the hopping between |$\ket{e}$ and |$\ket{f}$, $\Gamma\geq 0$ is the decay rate from |$\ket{e}$ to |$\ket{g}$ and $\hat{\Pi}=\ket{e}\bra{e}$ | denotes the projector over state |$\ket{e}$. The (shifted) spectrum $\varepsilon_{\pm}+i\frac{\Gamma}{2}=\pm\sqrt{J^{2}-\frac{\Gamma^{2}}{4}}$ is real for $J\geq\frac{\Gamma}{2}$, and imaginary for $J<\frac{\Gamma}{2}$. These two regimes respectively correspond to (passive) $\mathcal{PT}$ unbroken and broken symmetry phases Bender and Boettcher (1998).
@@
Figure 1: Purity evolution for different strengths of the noise, setting $\Gamma=1$, as given by (9) in a cross-section of the Bloch sphere. The stable steady states corresponding to parameters $(\Gamma/J,\gamma J)$, the hopping being increased in the range $J\in[10^{-2},10^{2}]$ (green line— See App. D for the analytical expressions).
@@
Figure 2: Spectral and steady-state phase diagrams: (a) dissipative gap $\Delta$ dictating the time-scale of convergence to the steady-state, (b) maximum imaginary part of the eigenvalues $\omega_{\nu}$, (c) $z$ and (d) $y$ components of the Bloch vector for the stable steady state $\hat{\rho}^{\textsc{s}}$. The $z$ coordinate (c) characterizes the three different phases to which the dynamics can stabilize. The dimensionless parameters determine the noise strength $\gamma J$ and the decay rate $\Gamma/J$. The transition to the noise-induced phase is at $\gamma^{*}=1/(2\Gamma)$ (black line). The transition from $\mathcal{PT}$ broken to the $\mathcal{PT}$ unbroken phase happens at $\Gamma/J=2$ (gray dashed line). In (b) the power laws $\gamma=1/(3\Gamma)$ and $\gamma=1/\Gamma$ are also shown (orange dotted lines).
@@
which physically corresponds to considering Gaussian fluctuations in the decay parameter $\Gamma$, with strength $\gamma$. This system can be experimentally realized with superconducting qubits Naghiloo et al. (2019), where noise is known to be present in the setup (cf. App. B.1), or in trapped ions Quinn et al. (2023), where noise may be easier to tune (cf. App. B.3). The master equation (4) for this system reads